OpenStudy (anonymous):

Can someone help? Find the maximum y-value on the graph of y= f(x). f(x) = -x^2 + 6x + 3

OpenStudy (anonymous):

If you graph it, the max value is at (3, 12), so the max y value is 12...

OpenStudy (anonymous):

thanks! @Amada How would I graph this? y = x^2 - 7

OpenStudy (anonymous):

I know the graph is downward because of -7

OpenStudy (anonymous):

You mean how to draw it, or how to graph it on a calculator? 

OpenStudy (anonymous):

Is what it should look like...

OpenStudy (anonymous):

Draw it, Basically the question is this: a)Sketch a graph of the equation. b)identify the vertex c)compare the graph of y=f(x) to the graph of y=x^2

OpenStudy (anonymous):

You plug in points for x to get y values and then you plot them...so at x=-2, y=-3, x= -1, y=-6, etc...

OpenStudy (anonymous):

The vertex point is a max or min, in this case the min is at x=0, y= -7

OpenStudy (anonymous):

Just plot both graphs for part c...by plugging in values for x to find the y

OpenStudy (anonymous):